Annie has quarters and dimes in her bank. There are eight less than four times as many dimes as quarters. There is $6.35 in her bank. How many dimes are in her bank?

Answer by wuwei96815(245)

The biggest problem with a word problem is interpreting the words and translating them into a math problem that can be sloved. Let's try.

Let x=the number of quarters. Then, 4x-8=the numbers of dimes.

The following equations shows the relationship between the number of coins and their total value:
$0.25(x) + $0.10(4x-8) = $6.35
.25x + .40x - .80 = 6.35
.65x = 6.35 + .80 = 7.15
x = 7.15/.65 = 11
4x-8 = 44-8 = 36